# Image is All You Need: Towards Efficient and Effective Large Language Model-Based Recommender Systems

### v1

@@ 3.5. Inference
4. Experiments
  • RQ3: How does each module (i.e., ILA and IRE) contribute to the effectiveness of I-LLMRec?
  • RQ4: How can we handle when item images are missing?
4.1. Experimental Setting
Datasets. For evaluation, we use four categories from the Amazon dataset (Ni et al., 2019): Sports, Grocery, Art, and Phone. These datasets contain the item titles, attributes (e.g., brand and category), detailed textual descriptions, and images representing items. Following a prior study (Wei et al., 2024), we filter out users and items with fewer than 5 interactions to ensure data quality. We summarize the data statistics in the Appendix F.
@@ 4.1. Experimental Setting
  • A-LLMRec (Kim et al., 2024a) express items by projecting the CF item embeddings into the LLM along with title, enabling effective recommendation of warm items.
  • TRSR (Zheng et al., 2024) is a representative work of the Description-based Representation approach, which summarizes full item descriptions to effectively express items while leveraging rich textual semantics.
@@
  • I-LLMRec+$\mathbf{D}$ (equiv. to I-LLMRec+TRSR) [^8] For each item in the user interaction, we append the description as ”Description: $\mathbf{D}_{i}$” after $\mathbf{P}_{i}$ within the prompt, i.e., $\mathbf{P}_{i}=\{\textrm{Title}:\textsf{ITEM\_TITLE},\textrm{Visual}\;\textrm{Representation}:[\textsf{VISUAL}],\textrm{Description}:\textbf{D}_{i}$ is a variant of I-LLMRec that represent items using both image features and text description.
@@
4.1.1. Implementation Details
We employed Sheared LLaMA 2.7B (Xia et al., 2023) as the LLM backbone. For the visual encoder $V$, we adopt SigLIP (Zhai et al., 2023), a variant of CLIP (Radford et al., 2021) with a sigmoid loss function, to extract visual features. Regarding experiments with different LLMs and visual encoders, please refer to Appendix G. We use a 2-layer MLP with ReLU activation function for the adaptor $M$. In IRE module, we employ SBERT (Reimers, 2019) as the textual encoder. To ensure consistency across the LLM backbone, we use the same Sheared LLaMA 2.7B for TALLRec, A-LLMRec, and TRSR. Since these models are limited to recommending items only from a narrower item pool (i.e., 1 or 20), we apply the same IRE module for recommendation, which incorporates image, CF, and textual features, allowing them to recommend items from a larger pool. This approach is acceptable since our main goal is to evaluate how effectively the LLM captures user preferences through item expression. For TRSR, we use GPT-4o (OpenAI, 2024) to summarize full item descriptions. For UniMP, we follow its original implementation, which utilizes OpenFlamingo-4B-Instruct (Awadalla et al., 2023), a pre-trained multimodal LLM. Regarding training configurations, we set the learning rate to 0.001 and the batch size to 32. All models are trained on an NVIDIA GeForce A6000 48GB GPU.
4.2. Performance Comparison (RQ1)
Table 1 shows the performance comparison between I-LLMRec and baselines across four datasets. The key observations are as follows: 1) LLM-based models generally outperform traditional CF models, highlighting the effectiveness of leveraging an LLM backbone for the recommender systems. 2) A-LLMRec and TRSR generally outperform TALLRec. It suggests that beyond the attributes, incorporating CF item embeddings or item descriptions further enhances the LLM’s ability to capture user preferences. However, we observe that A-LLMRec performs worse on the Phone dataset, and TRSR generally outperforms A-LLMRec, indicating that item descriptions are more effective than ID-based item embeddings in expressing item semantics. 3) I-LLMRec outperforms UniMP, an image-aware LLM-based model. Given that UniMP employs a multimodal LLM where images and texts are pre-aligned, it indicates the need for a more specialized alignment strategy tailored to the recommendation tasks, demonstrating the effectiveness of I-LLMRec. 4) Considering that TRSR includes attributes during summarization (See Appendix B), adding description information in TALLRec, which is equivalent to TRSR, improved performance, whereas descriptions in addition to images (i.e., I-LLMRec+$\mathbf{D}$) did not improve the performance of I-LLMRec. This suggests that while descriptions convey useful information, they are inherently limited by what is already present in images due to the significant information overlap between image and description pairs. Even if descriptions may include some text-specific information that images do not contain, the fact that performance remains competitive implies that such information has surprisingly little impact on the recommendations. This supports the idea that item images are sufficient to capture user preferences for recommendation.
4.3. Model Analysis (RQ2)
In this section, we analyze how I-LLMRec, which represents items using images, performs in recommendations compared to the natural language-based approach (TRSR, I-LLMRec+$\mathbf{D}$, and TALLRec) in terms of efficiency, effectiveness, and robustness.
Figure 4. Inference time (seconds/100 users) over the length of users’ item interaction sequence (i.e., $|\mathcal{S}_{u}|$).
4.3.1. Analysis on Efficiency.
To study the efficiency of I-LLMRec, we evaluate the inference time across different sizes of user’s item interaction sequences (i.e., $|\mathcal{S}_{u}|$). Specifically, we divided the users into groups according to $|\mathcal{S}_{u}|$, randomly selected 100 users per group, and measured their total inference time. As shown in Figure 4, we observe that TRSR and I-LLMRec+$\mathbf{D}$, which rely on lengthy descriptions, show a sharp increase in inference time as the length of textual descriptions within the prompt increases proportional to $|\mathcal{S}_{u}|$. On the other hand, I-LLMRec maintains consistently low inference time regardless of $|\mathcal{S}_{u}|$. This efficiency stems from the fact that even as $|\mathcal{S}_{u}|$ increases, only a minimal number of image tokens are added, keeping computational costs low. Furthermore, we observe that TALLRec, which is an attribute-based representation approach, shows a gradual increase in inference time and eventually surpasses I-LLMRec when $|\mathcal{S}_{u}|\geq 10$. We attribute this to the fact that processing natural language attributes scales in complexity faster than image processing required for I-LLMRec. In summary, I-LLMRec is more efficient than TRSR across all $\mathcal{S}_{u}$ and even more efficient than TALLRec when user interactions are high, demonstrating superior computational efficiency.
Figure 5. Performance of various LLM context window sizes.
4.3.2. Analysis on Effectiveness.
To gain a deeper understanding of the effectiveness of I-LLMRec, we examine how well it performs under a limited budget compared to the natural language-based approach. Specifically, we consider the LLM’s context window size (i.e., maximum input token length) as the budget, and vary it from 4,096 to 256 tokens. This allows us to analyze the ability of models in effectively capturing the user preferences given a limited budget. As shown in Figure 5, we observe that TRSR, I-LLMRec+$\mathbf{D}$, and I-LLMRec perform consistently when the context window size is over 1,024 since they can handle most interactions. However, as the context window size drops below 512, the performance of TRSR and I-LLMRec+TRSR declines sharply. This is because lengthy descriptions make it difficult for the LLM to fully process interactions with a limited context window size, thereby making it more challenging to understand user preferences. On the other hand, I-LLMRec maintains stable performance even at the context window size of 256 by representing item semantics using minimal image tokens. This enables the LLM to process full user interactions within the given budget and effectively capture user preferences. While TALLRec also remains stable at 256 due to its low token usage, it exhibits inferior performance because its attributes are not expressive enough. Overall, I-LLMRec effectively overcomes the context window size constraints by leveraging item images to preserve rich item semantics, demonstrating its effectiveness and practicality.
Figure 6. Performance of natural language-based approaches (TALLRec, TRSR and Full Description) and I-LLMRec on four datasets (Sports, Grocery, Phone and Art).
4.3.3. Analysis on Robustness
To evaluate the robustness of representing items using images instead of natural language as done in I-LLMRec, we compare its performance with various natural language-based approaches—TALLRec (Attribute), TRSR (Summarized description), and Full description—across multiple datasets. Figure 6 presents two key findings regarding the drawback of natural language-based approaches. 1) Information loss from summarization (Figure 6(a)): Full description outperforms TRSR, indicating that summarizing descriptions introduces information loss, where essential semantic details may be omitted, leading to lower performance. 2) Presence of Noise in Full Description (Figure 6(b)): On some datasets (i.e., Phone and Art), Full description rather performs worse than TRSR that relies on a summarized description. This indicates that the lengthy full description may contain noise and that the noise could be partially removed in the summarized version of the full description [^9] Indeed, we found that full descriptions often contain extraneous content, such as HTML tags introduced during data crawling.. Especially in Phone dataset, Full Description even underperforms TALLRec that relies only on simple attributes. Such different behaviors of natural language-based approaches across different datasets demonstrates that these approaches are sensitive to the given text, and thus impractical in reality. In contrast, I-LLMRec, which captures item semantics through images rather than text, consistently delivers strong performance across all datasets, demonstrating its robustness in handling variations in item descriptions.
Table 2. Ablation studies of I-LLMRec.
Row ILA IRE Sport Art
Image CF Text Hit@5 NDCG@5 Hit@5 NDCG@5
(a) 0.3953 0.3043 0.5040 0.3915
(b) 0.4316 0.3403 0.5564 0.4447
(c) 0.4075 0.3256 0.5502 0.4517
(d) 0.4491 0.3630 0.5795 0.4769
(e) 0.4570 0.3711 0.5883 0.4839

### v2

@@ 4.1 Experimental Setting
4.2 Performance Comparison (RQ1)
[…]
4.3 Model Analysis (RQ2)
Analysis on Efficiency.
Figure 3: Inference time (sec/100 users) over $|\mathcal{S}_{u}|$.
[…]
Figure 4: Performance of various LLM context window sizes.
[…]
Figure 5: Performance of natural language-based approaches (TALLRec, TRSR and Full Description) and I-LLMRec on four datasets.
Analysis on Robustness. To evaluate the robustness of representing items using images instead of natural language as done in I-LLMRec, we compare its performance with various natural language-based approaches—TALLRec (Attribute), TRSR (Summarized description), and Full description—across multiple datasets. Figure 5 presents two key findings regarding the drawback of natural language-based approaches. 1) Information loss from summarization (Figure 5(a)): Full description outperforms TRSR, indicating that summarizing descriptions introduces information loss, where essential semantic details may be omitted, leading to lower performance. 2) Presence of Noise in Full Description (Figure 5(b)): On some datasets (i.e., Phone and Art), Full description rather performs worse than TRSR that relies on a summarized description. This indicates that the lengthy full description may contain noise and that the noise could be partially removed in the summarized version of the full description [^10] Indeed, we found that full descriptions often contain extraneous content, such as HTML tags introduced during data crawling.. Such different behaviors of natural language-based approaches across different datasets demonstrates that these approaches are sensitive to the given text, and thus impractical in reality. In contrast, I-LLMRec, consistently delivers strong performance across all datasets, demonstrating its robustness in handling variations in item descriptions.
Table 2: Ablation studies of I-LLMRec.
Row RISA RERI Sport Art
Image CF Text Hit@5 NDCG@5 Hit@5 NDCG@5
(a) 0.3953 0.3043 0.5040 0.3915
(b) 0.4316 0.3403 0.5564 0.4447
(c) 0.4075 0.3256 0.5502 0.4517
(d) 0.4491 0.3630 0.5795 0.4769
(e) 0.4570 0.3711 0.5883 0.4839
Figure 6: Performance across item groups, where higher IDs indicate warmer item groups.
4.4 Ablation Studies (RQ3)
In Table 2, we conduct ablation studies to understand the effectiveness of each component in I-LLMRec. The variant without any additional components is trained solely using Equation 1 and recommends items based only on image features (row (a)). We make the following observations: 1) Effect of RISA: We observe that equipment with the RISA module significantly improves performance (row (a) vs. row (b)). This indicates that the RISA module effectively aligns visual features with the language space, enhancing the LLM’s ability to understand user preferences. 2) Effect of extending to multiple feature types in RERI: We observe that using two features together [^11] The inclusion of specific feature types is applied both in model training (Equation 2) yields better performance than using either the image or CF feature alone (row (b), (c) vs. (d)). Moreover, integrating all three feature types (i.e., Image, CF, and Text) further improves performance (row (d) vs. (e)). This demonstrates that extending multiple feature types enhances recommendation effectiveness beyond image features alone.
Effectiveness of Feature types across cold/warm items. To further analyze the impact of different feature types, we evaluate their performance on cold and warm item recommendations. Specifically, we follow (Liu et al., 2023b) by sorting the items in ascending order based on the frequency of interaction and dividing them into five equally sized groups, where a higher group ID indicates a warmer item group, Then, we compare the recommendation performance across these groups. As shown in Figure 6, we observe that using image features (blue bar) generally outperforms CF features (orange bar) for cold items (lower ID groups). However, as the item group transitions from cold to warm, CF’s effectiveness improves, eventually surpassing Image in Item Group 5 (Warm). This suggests that CF is particularly effective for recommending warm items. With these insights, combining Image and CF (green bar) helps mitigate CF’s weakness in recommending cold items, resulting in higher performance than CF alone in Item Group 1-2. At the same time, the CF features compensate for weakness of the image features in recommending warm items, leading to better performance than image alone in Item Group 4-5. Furthermore, adding textual features (red bar) further boosts performance across all Item Groups, exceeding the performance of Image+CF. This analysis highlights how different feature types complement each other to enhance recommendation.
5 Discussion
Rethinking Information Overlap. We note that the observation of information overlap between image and text features, as discussed in Section 1, relates to the challenges of redundancy in multimodal recommender systems (Zhou et al., 2023; Yang et al., 2025; Jeong et al., 2024). Here, we discuss how this information overlap has been perceived in previous studies and articulate how our viewpoint departs from these interpretations. Specifically, prior studies on multimodal recommender systems generally view this information overlap as redundant information that is seen as a barrier to improving performance when both modalities are used. Therefore, they generally aim to extract complementary information from each modality to improve performance. On the other hand, we present a fundamentally different viewpoint, asserting that this information overlap is in fact a crucial factor in addressing the trade-off between efficiency and effectiveness when representing items for LLMs in natural language. In other words, we view this information overlap as an advantage for LLM-based recommender systems rather than a barrier. Moreover, this perspective allows us to maintain robust performance in real-world cases where item images are absent, as detailed in Appendix E.8 (RQ4).
6 Related Work
LLM-based Recommendation. A myriad of studies have spurred the development of LLMs for recommender systems (Bao et al., 2023; Wang and Lim, 2023; Zhang et al., 2023; Liu et al., 2024b). Prior works adapt LLMs to recommendation by representing interaction history as item title sequences and enhancing performance with item attributes to enrich semantics. (Attribute-based Representation). Specifically, TALLRec (Bao et al., 2023) converts item IDs into their titles and captures user preference via LoRA (Hu et al., 2021) fine-tuning. TransRec (Lin et al., 2024b) improves the user preference understanding by including item titles and attributes in prompts. IDGenRec (Tan et al., 2024) incorporates item titles generated by an independent LLM for unique and meaningful semantics. More recently, to feed the richer item semantics to LLMs, a description of items have been utilized (Description-based Representation). Specifically, TRSR (Zheng et al., 2024b) utilizes a large LLM (i.e., LLaMA-30B-instruct (Touvron et al., 2023a)) to summarize item descriptions, and feeds them to a smaller LLM (i.e., LLaMA-2 7B (Touvron et al., 2023b)) to capture the condensed meaningful information, while ONCE (Liu et al., 2024b) adopts GPT-3.5 (Brown et al., 2020) for the summarizing task. Despite their notable success to capture the richer item semantics, they face a trade-off between efficiency and effectiveness when representing items in natural language, leading us to utilize image information to address this trade-off.
Please see Appendix F for additional related works (e.g., multimodal LLM-based recommendation).
7 Conclusion
In this work, we address the trade-off between efficiency and effectiveness when representing item semantics in natural language for an LLM-based recommender system. Based on the observation that there exists a significant information overlap between images and descriptions associated with items, we propose I-LLMRec, a novel method that leverages images to capture rich item semantics of lengthy descriptions with only a few tokens, thereby capturing user preferences efficiently and effectively. However, as the image and language space are not inherently aligned, we propose the Recommendation-oriented Image-LLM Semantic Alignment (RISA) module, which effectively bridges the gap between these spaces. Furthermore, we propose the REtrieval-based Recommendation via Image features (RERI) module, a retrieval-based recommendation approach, to enhance both the reliability and efficiency of the recommender systems. Our extensive experiments demonstrate that I-LLMRec outperforms natural language-based approaches in terms of both efficiency and effectiveness. Regarding the limitation and future work, please refer to Appendix G
Acknowledgement
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-II220157, Robust, Fair, Extensible Data-Centric Continual Learning) as well as another IITP grant funded by the Korea govern- ment(MSIT) (RS-2022-II220077, Reasoning, and Inference from Heterogeneous Data).
Ethics Statement
In accordance with the ICLR Code of Ethics, to the best of our knowledge, we have not encountered any ethical concerns in the course of this work. Furthermore, all datasets and pre-trained models used in our experiments are publicly available.
Supplementary Material
- Token-Efficient Item Representation via
Images for LLM Recommender Systems -
Figure 7: Prompt for summarizing descriptions. The red-colored text denotes the actual metadata for the item.
Appendix A Prompt for Summarization
Figure 7 shows the prompt used for description summarization. Following TRSR Zheng et al. (2024b), we incorporate attribute information along with the description to generate a concise summary.
Appendix B Recommendation Protocol
In this section, we describe a comprehensive explanation of the recommendation protocol, which is consistently applied to Attribute-based Representation, Description-based Representation (i.e., Full Description and Summarized Description), and I-LLMRec as introduced in Section 1. Note that this recommendation protocol follows a retrieval-based recommendation approach similar to the RERI module discussed in Section 3.3. Therefore, we recommend reading Section 3.3 first for a deeper understanding. At a high level, we modify the way items are represented by replacing the visual representation with either an attribute-based or description-based representation within $\mathbf{P}_{i}$, followed by a recommending process similar to RERI module.
Specifically, we first craft the prompt, which includes the task prompt (as shown in Figure 2) and the semantics of user-interacted items in a sequence. Here, the item semantics depend on the item representation approach, which plays a crucial role in capturing user preferences. Then, we append a learnable token $\left[\texttt{REC}\right]$ at the end, allowing this token to aggregate the semantics of items the user has interacted with, thereby capturing user preferences. To obtain the corresponding representation, we use the last hidden state of $\left[\texttt{REC}\right]$ token, which serves as the LLM-guided user representation. Using this representation, we retrieve relevant items by comparing with items’ visual, collaborative filtering (CF), and textual features. More precisely, for each feature type, we compute affinity scores by performing a dot-product between the LLM-guided user representation and item features. Based on the summation of affinity score across feature types for all items, we rank the items and recommend the Top-$k$ items with the highest scores.
Figure 8: Cosine similarity between item image-description pairs in Amazon Phone and Grocery datasets and image-caption pairs in the COCO dataset using CLIP Radford et al. (2021).
Figure 9: Sigmoid-based similarity (ranging from 0 to 1) between item image-description pairs across four datasets (Amazon Sport, Art, Grocery, and Phone datasets) using SigLIP Zhai et al. (2023).
Appendix C Consistent Observation of Information Overlap
We further investigate the information overlap between image and description associated with items, which was explored in Section 1, by expanding these experiments to additional datasets and leveraging another CLIP-based model. 1) Expanding to additional datasets: Building upon our previous comparative experiments in Amazon Sport and Art—where we used CLIP Radford et al. (2021) to measure cosine similarity in Section 1—we extend our analysis to the Amazon Grocery and Phone datasets. As shown in Figure 8, the average similarity for Amazon Phone and Grocery is approximately 0.31, which is higher than 0.26 average similarity observed in well-aligned COCO image-caption pairs. This consistent observation across multiple datasets further supports the information overlap between item images and descriptions. 2) Leveraging another CLIP model: To further explore the information overlap using a different CLIP variant, we select SigLIP Zhai et al. (2023), a variant of CLIP with sigmoid loss function, to compute the similarity scores ranging from 0 to 1. Specifically, when we apply SigLIP to measure similarity between item image-description pairs across all four datasets (Amazon Sport, Art, Grocery, and Phone), we observed that, as show in Figure 9, the majority of items exhibit similarity scores close to 1, further demonstrating the significant information overlap between their image and description pairs.
Appendix D Proposed Method: I-LLMRec
Figure 10: Prompt template for RISA module. The red-colored text denotes the actual information of an item.
D.1 RISA Prompt Templates
Figure 10 shows the prompt template for RISA module of  I-LLMRec. As described in Section 3.2, we utilize the prompt templates to align the visual feature with the language space of LLMs by training the adaptor network $M$.
D.2 Discussion of LLM-finetuning for I-LLMRec
A further benefit of I-LLMRec is that LLM fine-tuning is not mandatory, unlike the existing LLM-based recommender systems (Lin et al., 2024b; Bao et al., 2023; Tan et al., 2024; Zheng et al., 2024b; Liu et al., 2024b) that generally rely on costly fine-tuning of the LLM. Instead, we harness the LLM’s intrinsic parametric knowledge to understand the user preferences. Throughout this paper, we opt to keep the LLM frozen to preserve the LLM’s intrinsic knowledge and improve the training efficiency.
Table 3: Performance comparison when LLM backbone of  I-LLMRec is fine-tuned using LoRA Hu et al. (2021).
Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
LLM Frozen 0.4570 0.3711 0.5883 0.4839
LLM fine-tuning with LoRA 0.4633 0.3772 0.5958 0.4949
Potential Improvement through LLM fine-tuning Nevertheless, we explore potential performance gains from fine-tuning the LLM, a common approach in existing LLM-based recommender systems Lin et al. (2024b); Bao et al. (2023); Tan et al. (2024); Zheng et al. (2024b); Liu et al. (2024b). However, since fully fine-tuning an LLM demands substantial computational resources, we adopt the parameter-efficient fine-tuning strategy, LoRA Hu et al. (2021), which introduces a small number of learnable rank matrices into each LLM’s transformer layer. As shown in Table 3, we observed the performance improvement through fine-tuning, indicating that there is room for further improvement within I-LLMRec. However, considering the computational cost, we opt to freeze the LLM.
D.3 Discussion and Comparison with Semantic ID and Textual Identifier Approaches
Discussion with Semantic ID Approaches. Recently, generative recommender systems with semantic ID methods (Zheng et al., 2024a; Lin et al., 2025; Rajput et al., 2023; Kim et al., 2026; Hou et al., 2025) has been emerged. Here, we discuss the differences between our approach and semantic ID approach, and compare the efficiency and performance.
While both approaches share the goal of efficiently representing items within the LLM’s capacity, there is a fundamental difference in how they convey the semantics of item descriptions to the LLM. The Semantic ID approach addresses the challenges of randomly initialized item tokens—which inherently fail to capture semantic relations between items (Liu et al., 2025)—by training a vector-quantized model (Zeghidour et al., 2021) that maps item descriptions to discrete codes. These descriptions serve as semantic intermediaries, aiding in the capture of semantic relationships between items (i.e., if two items share similar descriptions, they typically have the same codes), and allowing the LLM to implicitly capture the semantics of the descriptions. On the other hand, our approach explicitly encodes the semantics of each item’s description using images, which provide a richer and more direct representation for the LLM.
Regarding token efficiency, Semantic ID approaches (Zheng et al., 2024a; Lin et al., 2025) generally require four tokens (codes) per item, whereas our approach uses a single image token that captures the semantics of the description. Importantly, Semantic ID approaches essentially require multiple discrete tokens to express cross-item semantic relationships, while our approach conveys the semantics of the description with just one image token. Beyond token efficiency, we also compare performance with a state-of-the-art semantic ID method, LC-Rec (Zheng et al., 2024a), under a fair evaluation setting using the same LLM backbone (Xia et al., 2023) on the Sport and Art datasets with full item candidates. I-LLMRec achieves Hit@5 scores of 16.28 on the Sport dataset and 21.80 on the Art dataset, compared to 15.86 and 20.41 for LC-Rec, respectively. These results indicate that I-LLMRec not only offers more efficient item representation but also more effective in representing item semantics to the LLMs.
Discussion with Textual Item Identifier. While Semantic ID approaches rely on latent codes (tokens) for item representation, the textual identifier methods (e.g., IDGenRec (Tan et al., 2024)) generally generate a paraphrased version of an item’s title, derived from its textual description, to serve as its identifier. Similar to Semantic ID approaches discussed earlier, this approach also shares the goal of efficiently representing items. However, the textual identifiers inevitably lose the semantics of descriptions since they compress the full item description into a shortened title-like phrase. As a result, the state-of-the-art textual identifier method, IDGenRec, achieves only 15.11 Hit@5 on the Sport dataset and 17.97 on the Art dataset under the full item candidate setting—performance that is inferior to I-LLMRec (16.28 and 21.80, respectively) and still below that of the semantic ID method LC-Rec.
Regarding token efficiency, IDGenRec compresses roughly 209 tokens of description into an average of 10.9 tokens per item. On the other hand, I-LLMRec uses a single visual token that directly encodes the item’s semantic content. This leads to significantly lower computational cost during both training and inference, while still retaining richer semantic information.
Appendix E Experiments
Table 4: Data statistics after preprocessing.
Dataset # Users # Items # Interactions Avg. Len
Sport 24,665 9,884 179,491 7.28
Grocery 84,896 25,632 739,628 8.71
Art 14,986 5,801 121,917 8.13
Phone 59,307 19,671 403,194 6.80
E.1 Data Statistics
E.2 Implementation Details
We employed Sheared LLaMA 2.7B (Xia et al., 2023) as the LLM backbone. For the visual encoder $V$, we adopt SigLIP (Zhai et al., 2023), a variant of CLIP (Radford et al., 2021) with a sigmoid loss function, to extract visual features. Regarding experiments with different LLMs and visual encoders, please refer to Section E.12. We use a 2-layer MLP with ReLU activation function for the adaptor $M$ and the intermediate dimension is 512. In RERI module, we employ SBERT (Reimers, 2019) as the textual encoder. To ensure consistency across the LLM backbone, we use the same Sheared LLaMA 2.7B for TALLRec, A-LLMRec, and TRSR. Since these models are limited to recommending items only from a narrower item pool (i.e., 1 or 20), we apply the same RERI module for recommendation, which incorporates image, CF, and textual features, allowing them to recommend items from a larger pool. This approach is acceptable since our main goal is to evaluate how effectively the LLM captures user preferences through item expression. For TRSR, we use GPT-4o (OpenAI, 2024) to summarize full item descriptions. For UniMP, we follow its original implementation, which utilizes OpenFlamingo-4B-Instruct (Awadalla et al., 2023), a pre-trained multimodal LLM. Regarding training configurations, we set the learning rate to 0.001 and the batch size to 32. All models are trained on an NVIDIA GeForce A6000 48GB GPU.
E.3 Baselines
@@
  • A-LLMRec (Kim et al., 2024b) express items by projecting the CF item embeddings into the LLM along with title, enabling effective recommendation of warm items.
  • TRSR (Zheng et al., 2024b) is a representative work of the Description-based Representation approach, which summarizes full item descriptions to effectively express items while leveraging rich textual semantics.
@@
[…]
@@
Table 5: Performance results on three additional Amazon datasets (Automotive, Video, and Clothing) and the H&M Fashion dataset, with cosine similarity (Cosine Sim.) between item images and descriptions measured using CLIP (Radford et al., 2021) is measured.
Model
Automotive
(Cosine Sim.: 0.3086)
Video
(Cosine Sim.: 0.2801)
Clothing
(Cosine Sim.: 0.3107)
H&M Fashion
(Cosine Sim.: 0.2847)
Hit@5 NDCG@5 Hit@5 NDCG@5 Hit@5 NDCG@5 Hit@5 NDCG@5
SASRec 0.4027 0.3078 0.6045 0.4713 0.4981 0.4690 0.4918 0.3885
TALLRec 0.4203 0.3107 0.6040 0.4521 0.4388 0.3660 0.3959 0.2827
I-LLMRec 0.4515 0.3379 0.6413 0.4915 0.5469 0.4927 0.5321 0.3949
E.4 Analysis on Other Datasets
Behavior on visual-centric datasets. We evaluate on two visual-centric datasets (Amazon Clothing and H&M Fashion datasets), where users’ choices are likely influenced by visual attributes, and two less visual-centric datasets (Amazon Automotive and Video) to examine how I-LLMRec performs in visual-centric datasets compared to baselines, including the collaborative filtering model (SASRec(Kang and McAuley, 2018)) and natural language-based LLM approach (TALLRec (Bao et al., 2023)). As shown in Table 5, we observe that TALLRec performs on par with SASRec on the less visual-centric datasets, while showing a significant performance drop on visual-centric ones. It indicates that natural language alone is insufficient to represent items for LLMs in vision-driven contexts, making it challenging to capture user preferences driven by visual attributes. I-LLMRec, however, effectively leverages image information, yielding significantly higher performance than TALLRec, especially on visual-centric datasets. These results highlight the particular effectiveness of I-LLMRec in vision-driven contexts, while its solid performance on Automotive and Video further demonstrate its general effectiveness.
Generalization of the observation. To further demonstrate the information overlap between item images and descriptions discussed in Section 1, we measured the cosine similarity between images and descriptions on three additional Amazon datasets (Automotive, Video, and Clothing) as well as the H&M Fashion dataset (Ling et al., 2022) from another domain. As shown in Table 5, the high cosine similarity (Cosine Sim.) reported under each dataset consistently reveals a strong semantic overlap between images and descriptions, alongside superior performance of I-LLMRec over the baselines. While we could not cover every real-world dataset, the consistent results across eight datasets—including the four in the main paper—demonstrate the general applicability of information overlap, suggesting that I-LLMRec can be broadly applied.
Table 6: Performance on MicroLens dataset.
Hit@5 NDCG@5
SASRec 0.4567 0.3413
TALLRec 0.4459 0.3237
I-LLMRec 0.4854 0.3520
Generalization to the Video Domain. We extended our evaluation to the micro-video recommendation domain by using the MicroLens (Ni et al., 2025) dataset and compared the performance with SASRec and TALLRec baselines. For visual features v{v}, we average the image features of sampled frames provided by the MicroLens dataset. The results, as shown in the Table 6, show that TALLRec performs similarly to a traditional collaborative filtering model (SASRec), highlighting that user preferences based solely on titles are insufficient. On the other hand, I-LLMRec significantly outperforms TALLRec, demonstrating the generalization of I-LLMRec to the video domain.
Table 7: Performance on Goodbooks dataset.
Hit@5 NDCG@5
SASRec 0.4475 0.2956
TALLRec 0.4752 0.3266
TRSR 0.5100 0.3551
I-LLMRec 0.4870 0.3372
Behavior on Visually-Weak Domains. To explore the behavior of I-LLMRec on visually weak domains, we conducted additional experiments on the Goodbooks [^13] https://www.kaggle.com/datasets/zygmunt/goodbooks-10k dataset with multiple baselines (SASRec, TALLRec, and TRSR). As shown in Table 7, the description-based model TRSR outperforms I-LLMRec on this dataset. This aligns with intuition: in visually weak domains, visual features carry limited information for capturing user preferences, and textual descriptions naturally become more informative. However, we highlight two important points. First, while long textual descriptions can help in fully text-driven domains like Goodbooks, relying on them is often impractical due to significant computational cost. For example, TRSR incurs 2.8× longer inference time and 3.7× longer training time than I-LLMRec on Goodbooks. Furthermore, although we can leverage descriptions in such domains due to their low image–description cosine similarity (0.16 compared to over 0.29 in other datasets), these domains are limited in scope. Second, across the eight datasets used in the paper—including Sport and Phone, which can also be considered visually weak—we consistently observe the high degree of semantic overlap between item images and descriptions, and I-LLMRec performs effectively, demonstrating that I-LLMRec generalizes beyond purely visually rich domains (e.g., Fashion). In summary, the efficiency and effectiveness of I-LLMRec across diverse domains make it broadly more suitable for real-world recommendation scenarios.
E.5 Regarding the Precomputation of Visual Features
Here, we clarify how visual feature extraction is handled in our efficiency evaluation. Following prior multimodal recommendation studies (Xv et al., 2024; Xu et al., 2025; Lin et al., 2024a), we pre-compute and cache all item visual features before both training and inference. Thus, the wall-clock times in Figure 3 of the main paper exclude visual feature extraction cost. However, we argue that this setup aligns with real-world recommendation scenarios: once an item’s visual features are extracted offline, the same representation can be reused for every user who interacts with that item, during both training and inference. Prior work (Xv et al., 2024; Xu et al., 2025) adopts the same convention when evaluating efficiency. Furthermore, it is important to note that caching these features is inexpensive—only 3.4KB per item, making the memory requirement extremely lightweight.
To conduct a thorough comparison of inference time with the description-based baseline (TRSR), we measure the inference time including visual feature extraction. On the Sports dataset, TRSR takes approximately 47 minutes for recommendation inference, while  I-LLMRec requires 15 minutes for recommendation inference and an additional 5 minutes for visual feature extraction, totaling around 20 minutes. Even with this setup,  I-LLMRec is 2.4 times more efficient.
E.6 Effectiveness of Image Features on Unseen Items
Extending the analysis of cold-start items in Section 4.4 of the main paper, to further demonstrate the effectiveness of image features for cold-start items, we evaluate the Sports and Art datasets by selecting users whose test items do not appear in the training set (i.e., extreme cold-start case), resulting in 302 users in Sports and 29 in Art. In this setting, using only CF features in the RERI module leads to near-zero performance, since unseen items retain untrained initial embeddings. On the other hand, image features achieve a Hit@5 of 0.3245 on Sports and 0.0689 on Art, demonstrating their effectiveness in recommending extreme cold start items.